OpenStudy (anonymous):

find the second derivative of x^2-xy=1-y^2. express as a function of x and y only

OpenStudy (anonymous):

what problem are u having?

OpenStudy (anonymous):

i dont know how to do it

OpenStudy (anonymous):

if u know how to differentiate,it is a pretty simple question then

OpenStudy (anonymous):

i never did this with xy and 2 x

OpenStudy (callisto):

Derivative with respect to x or y?

OpenStudy (anonymous):

yes can you show me a step by step of how to start it

OpenStudy (callisto):

That is a question for you.. Should we differentiate it with respect to x or y?

OpenStudy (anonymous):

well it says express as a function of x and y , so i am not really sure

OpenStudy (callisto):

With respect to x then... Usual practice :| e.g: \[\frac{d}{dx}(xy)=y\frac{d}{dx}x + x\frac{d}{dx}y=...?\]

OpenStudy (anonymous):

so should i brind all the numbers over to one side and do the first derivative then the second derivative

OpenStudy (callisto):

Better to do it step by step. Find the first derivative first.

OpenStudy (callisto):

Do you know implicit differentiation?

OpenStudy (anonymous):

um i am bad with the names but probably not but do i wright the problem as 2x=-2x

OpenStudy (anonymous):

because x^2 become 2x and -xy turn to one and 1 goes away and -y^2 truns into -2y

OpenStudy (callisto):

-xy won't turn to one and and -y^2 is not -2y

OpenStudy (anonymous):

then can you show me the way the formula should look next so i can see what to do

OpenStudy (aravindg):

Hi there! Welcome to Openstudy :) If you have any doubts on how the site works ,feel free to ask me or any other Ambassador with purple A near them .Have great time here . Check out our COC : http://openstudy.com/code-of-conduct

OpenStudy (callisto):

x^2-xy=1-y^2 Differentiate both sides with respect to x \[\frac{d}{dx}(x^2-xy)=\frac{d}{dx}(1-y^2)\]\[\frac{d}{dx}(x^2)-\frac{d}{dx}(xy)=\frac{d}{dx}(1)-\frac{d}{dx}(y^2)\]\[2x-[y\frac{d}{dx}(x)+x\frac{d}{dx}(y)]=-2y\frac{dy}{dx}\]\[2x-y-x\frac{dy}{dx}=-2y\frac{dy}{dx}\]\[2x-y=-2y\frac{dy}{dx}+x\frac{dy}{dx}\]\[2x-y=(x-2y)\frac{dy}{dx}\]\[\frac{dy}{dx}=\frac{2x-y}{(x-2y)}\]That's what you'll get for the first derivative.

OpenStudy (anonymous):

thank you that more clear i have to see it to learn it

OpenStudy (anonymous):

so i have to now find the second derivative right

OpenStudy (anonymous):

what do i do next?

OpenStudy (callisto):

You need to know implicit differentiation. <Example 1> Make use product rule. \[\frac{d}{dx}(xy) = x\frac{d}{dx}(y)+y\frac{d}{dx}(x)=x\frac{dy}{dx}+y\] <Example 2> Make use of power rule and chain rule. \[\frac{d}{dx}(y^n)=(n)y^{n-1}\frac{dy}{dx}\] Yes, you have to differentiate the equation given twice.

OpenStudy (anonymous):

of so for the 2 derivative do you get 3(x-y)/(x-2y)^2

OpenStudy (callisto):

No.. You need to differentiate it implicitly.

OpenStudy (anonymous):

o... can u show me?

OpenStudy (callisto):

You need to read how I get the first derivative first.

OpenStudy (callisto):

If you don't understand, ask. Otherwise, you can't get the second derivative correct.

OpenStudy (anonymous):

i thought that the way to get the first derivative is the same way to get the socond serivative its just done 2

OpenStudy (callisto):

Yes. but how do you get the first derivative?

OpenStudy (anonymous):

dident you just show me you take the numver like 3^4 and do 4*3= 12^3 per number.

OpenStudy (callisto):

I showed you.. But you don't seem to get it.

terenzreignz (terenzreignz):

WELCOME TO IMPLICIT DIFFERENTIATION 101 ! FEATURING TONIGHT'S GUEST @romio So, Romio, it seems you're having trouble differentiating implicitly. Let's have recap. The following questions require you to differentiate with respect to x? Ready? Here we go. What is... \[\huge \frac{d}{dx}x^4\]

OpenStudy (anonymous):

4x^3

terenzreignz (terenzreignz):

The crowd is going WILD :D Nicely done, @romio Differentiating implicitly involves a twist... another variable, usually y, is involved So, for the second question, what do you think is... \[\huge \frac{d}{dx}y^4\]? Remember, implicit differentiation assumes y is a function of x. So... your verdict, @romio ?

OpenStudy (anonymous):

i have now (2x-y)/(x-2y) as the first derivative and apparently i tried the second derivative and i got that wrong

OpenStudy (anonymous):

i need someone to show me a step by step for the second derivative

terenzreignz (terenzreignz):

Well, if you must, and you're sure your first derivative is correct, just redo the steps, differentiating with respect to x assuming y is some function of x. Whatever you did in getting the first derivative, it's the same as in getting the second.

OpenStudy (anonymous):

i know i did it, but Callisto said my anwer was wrong so i dont know what i did wrong

terenzreignz (terenzreignz):

Well, the only way to find out, is by redoing the question, again :D Don't worry... we are under the ever-watchful eye of @Callisto ;) Let's start with your equation... \[\large x^2-xy=1-y^2\] Now, your challenge, should you choose to accept it, is to differentiate both sides of the equation with respect to x. Are you up for it? :) \[\huge \frac{d}{dx}(x^2-xy)=\frac{d}{dx}(1-y^2)\]

terenzreignz (terenzreignz):

You'd do well to remember that differentiation distributes over addition and subtraction. Good hunting :D

OpenStudy (anonymous):

i got (2x-y)/(x-2y) for my first derivative

terenzreignz (terenzreignz):

Yeah. But Callisto said it's wrong, didn't she? Well then, let's redo the question. I'm not saying I agree with her or disagree, but the only way to find out is to do it, step by step, as you said. Don't be discouraged, just pause, do it slowly, and eventually, it'll be second nature. Now, about that question \[\huge \frac{d}{dx}(x^2-xy)=\frac{d}{dx}(1-y^2)\]

OpenStudy (anonymous):

no the second derivative the first was wright

terenzreignz (terenzreignz):

Was it now? Can you tell me what \[\Large \frac{d}{dx}y^4\]is?

OpenStudy (anonymous):

4y^3

OpenStudy (anonymous):

please can you show me to to turn my first derivative into a second derivative i did it like 20 times and i am doing something wrong

terenzreignz (terenzreignz):

I see... what the problem is :) Indeed \[\Large 4y^3=\frac{d}{dy}y^4 \]but \[\Large \frac{d}{dx}y^4=?\]

terenzreignz (terenzreignz):

The thing is, we are differentiating along x, and y is a FUNCTION of x, so a good thing to remember, is to differentiate y AS YOU WOULD DO if it were x, however, when you're finished, you append dy/dx (or y') to it, due to the chain rule... \[\Large \frac{d}{dx}y^4=4y^3\frac{dy}{dx}=4y^3y'\] I prefer the former, the one involving dy/dx, it's more... in-your-face :D

OpenStudy (anonymous):

ok so i took my first derivative and i did ((x-2y)(2-1)-[(2x-y)(1-2)])/(x-2y)^2 is this right my first derivatinv was (2x-y)/(x-2y)

OpenStudy (anonymous):

then i worked it out but am i on the right track

terenzreignz (terenzreignz):

Well, it seems like you've gotten it.

OpenStudy (anonymous):

yea i know but my answer was wrong can you do it and tell me what you got because i must be doing some kind of multiplication error or something like as brakedown from where i showed you

terenzreignz (terenzreignz):

Well, your first derivative is correct... how did you go about getting your second derivative?

OpenStudy (anonymous):

i used this formula ((x-2y)(2-1)-[(2x-y)(1-2)])/(x-2y)^2

OpenStudy (anonymous):

what is the answer T_T

OpenStudy (callisto):

The formula you used is not correct.

OpenStudy (anonymous):

o? what, why not i thought it was?.... ok well can you tell me what the correct formula to use for this is then?

terenzreignz (terenzreignz):

What formula? You just differentiate, don't you?

OpenStudy (callisto):

I've explained how to do it twice. Please refer to the above comments.

OpenStudy (anonymous):

so is the answer -3y/(x-2y)^2 for the second derivative and if that is the answer is that all i have to do

terenzreignz (terenzreignz):

Also, quotient rule... \[\Large \frac{d}{dx}\frac{f(x)}{g(x)}=\frac{f'(x)g(x)-f(x)g'(x)}{[g(x)]^2}\]

OpenStudy (anonymous):

it says express a function of x and y only

terenzreignz (terenzreignz):

@romio Wolfram Alpha? Two things: One, you can't do that on exams... Two, that gives you what's called the PARTIAL derivative of (2x-y)/(x-2y) \[\huge \frac{\partial}{\partial x}\frac{2x-y}{x-2y}=\frac{-3y}{(x-2y)^2}\]

OpenStudy (anonymous):

so is my answer wrong

terenzreignz (terenzreignz):

Yep... as far as the second derivative goes. Look, you can't escape learning how to differentiate implicitly, so just face the music ;)

OpenStudy (anonymous):

ok so can you tell me what the answer is so i can figure out what i am doing wrong

terenzreignz (terenzreignz):

Very poor choice of words when you know a Moderator is watching :D The Code of Conduct encourages you to arrive at the answer yourself, with, of course, some help from the users :)

OpenStudy (callisto):

Hmm... Do you know where you get it wrong in the first derivative, after I have given you the answer?

OpenStudy (anonymous):

partly, it will also be a good way to check if i did it wright because i dont know the answer so i wont be able to know if i am doing it right or wrong.

terenzreignz (terenzreignz):

Yeah, but there's a better way... actually doing it.... you know, taking \[\huge \frac{dy}{dx}=\frac{2x-y}{x-2y}\]and differtiating both sides? :D C'mon, I know you can do it ;)

terenzreignz (terenzreignz):

@romio has left?